A Cannon is fired with a muzzle velocity of 300 m/s, at an angle of 60 degrees. Find the

range, time of flight, maximum height attained by the projectile.


1
Expert's answer
2021-10-10T16:04:09-0400

The range, time of flight and maximum height are given as follows respectively (see https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/)


"R = \dfrac{u^2\sin2\theta}{g} \\
T = \dfrac{2u\sin\theta}{g}\\
h = \dfrac{(u\sin\theta)^2}{2g}"


where "u = 300m\/s" is the initial speed, "\theta =60\degree" is the launch angle, "g = 9.81m\/s^2" is the gravitational acceleration.

Substituting the values, obtain:


"R = \dfrac{300^2\cdot \sin(2\cdot 60\degree)}{9.81}\approx 7945m\\
T = \dfrac{2\cdot 300\cdot \sin60\degree}{9.81}\approx 53.0s\\
h = \dfrac{(300\cdot \sin60\degree)^2}{2\cdot 9.81} \approx 3440m"

Answer. 7945m, 53s, 3440m.